Predetermined Factory Overhead Rate Poehling Medical Center has a single operating room that is used by local physicians to perform surgical procedures. The cost of using the operating room is accumulated by each patient procedure and includes the direct materials costs (drugs and medical devices), physician surgical time, and operating room overhead. On January 1 of the current year, the annual operating room overhead is estimated to be:

Disposable supplies $341,600 Depreciation expense 61,600 Utilities 35,800 Nurse salaries 513,000 Technician wages 168,000 Total operating room overhead $1,120,000 The overhead costs will be assigned to procedures, based on the number of surgical room hours. Poehling Medical Center expects to use the operating room an average of eight hours per day, seven days per week. In addition, the operating room will be shut down two weeks per year for general repairs. a. Determine the predetermined operating room overhead rate for the year. $ per hour b. Bill Harris had a 8-hour procedure on January 22.

How much operating room overhead would be charged to his procedure, using the rate determined in part (a)? $ c. During January, the operating room was used 194 hours. The actual overhead costs incurred for January were $75,900. Determine the overapplied operating overhead or underapplied operating overhead for the period. Enter your answer as a positive number. $
Answer #1

Solution a:

Nos of estimated surgical hours = 8*7*50 = 2800 hours

Overhead rate = Estimated overhead / Estimated surgical hours = $1,120,000 / 2800 = $400 per hour

Solution b:

Operating room overhead to be charged to Bill harris = 8*$400 = $3,200

Solution c:

Operating room overhead applied in January = 194*$400 = $77,600

Actual operating room overhead for January = $75,900

Overapplied operating overhead = $77,600 - $75,900 = $1,700